# bibliography.bib

### v3

@@ 4 Gibbs sampler for a variance component model
We are given the yield of dyestuff from 6 batches of 5 yield measurements in grams (ie. $I=6$ and $J=5$). This data was published in Davies [dyestuff]. The data can be structured as a variance component model where $\theta_{i}$ represents the mean yield for batch $i$ and the priors are defined as follows,
$V\sim IG(6,6)\hskip 28.45274ptW\sim IG(6,6)\hskip 28.45274pt\mu\sim N(1600,1600)$
We want to find the estimated upper bound on the Wasserstein distance and total variation distance for the Gibbs sampler on the variance component model fitted to this data. Using the notation from Lemma 4.1, we set the initial values of our Markov chain to $x_{0}=(900,\ldots,900,100,200,900)$ and $X^{\prime}_{0}\sim\nu$ where $f_{\nu}$ is defined as in equation 7.
Applying R simulation and lemma 4.1, we estimate $K\leq 0.000489$. The following holds almost surely,
$W_{||\cdot||_{1},1}(\mathcal{L}(X_{n}),\pi)\leq 0.000489\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}||X_{n,i}-X^{\prime}_{n,i}||_{1}$
Using CRN technique, we simulated $||X_{n,i}-X^{\prime}_{n,i}||_{1}$, $300$ times ($M=300$) over 10,000 iterations ($N=10,000$). Figure 3 graphs the upper bound on the $L^{1}$ Wasserstein distance between $X_{n}$ and the corresponding stationary distribution. At iteration $9590$, the bound on the Wasserstein distance first hits below 0.01 and $W_{1}(\mathcal{L}(X_{9590}),\pi)\leq 0.009893$.
Figure 2: The thick black line represents the upper bound on the Wasserstein distance on the Gibbs sampler for a variance component model (Real Data example 4.1). The estimate uses the mean of 300 simulations with initial value $x_{0}=(900,\ldots,900,100,200,900)$. The light blue area represents the quantiles. The vertical axis is a log scale.
Real Data Example 4.2.
We use the same dataset as in Real Data example 4.1 with hyperparameters defined in Table 1.
We are going to compare our convergence bounds to the results from [simconv]. [simconv] generates an upper bound on the total variation distance between a Markov chain and its corresponding stationary distribution using a modified version of the popular DnM condition stated in [rose] (see Theorem 12 of [rose] and Proposition 1 of [simconv]) and estimates the required DnM parameters ($\Lambda,\lambda,\epsilon$) using simulation. The hyperparameters for the DnM bound are defined in Table 1.
@@
From [simconv] 0.5 1.0 0.0 0.0 0.0 $10^{12}$
This paper 0.5 1.00001 0.00001 1.00001 0.0 $10^{12}$
Table 1: Comparison of the hyperparameters used in Real Data Example 4.2 and [simconv]. The hyperparameters are slightly different to take into account limitations from Lemma 4.1, most notably that $b_{1},b_{2}>1$. The hyperparameters follow the same notation as in equation 6.
Proposition 24 of [wasmeth] will be used to bound the total variation of the variance component model given the Wasserstein distance. Note that in [wasmeth], the prior on $\mu\sim 1$ is flat. Since in this example $\mu\sim N(0,10^{12})$, we assume that they are the same.
Proposition 4.2 (Proposition 24 of [wasmeth]).
@@
$||\mathcal{L}(X_{n})-\pi||_{TV}\leq cJ^{3/2}IW_{||\cdot||_{2},p}(\mathcal{L}(X_{n-1}),\pi)$ (9)
where $c=c_{1}+c_{2}$ and
@@
We bound the whole chain in Proposition 4.2 $(\vec{\theta_{n}},V_{n},W_{n},\mu_{n})$ rather than the marginal chain $(\vec{\theta_{n}},\mu_{n})$ as presented in Proposition 24 of [wasmeth] due to de-initialization properties [roberts:rosenthal:2001]. See section 5.1 of [wasmeth] for more details.
Applying R simulation and lemma 4.1, we calculate $K=0.1228$. The following holds almost surely,
$W_{||\cdot||_{1},1}(\mathcal{L}(X_{n}),\pi)\leq 0.1228\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}||X_{n,i}-X^{\prime}_{n,i}||_{1}$
@@
Using Proposition 4.2 notation, $cJ^{3/2}I=563.178$ and so an upper bound on the total variation distance between the Markov chain and its corresponding stationary distribution is as follows,
$||\mathcal{L}(X_{n})-\pi||_{TV}$ $\leq 69.158\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}||X_{n,i}-X^{\prime}_{n,i}||_{2}$
Using CRN technique, we simulated $||X_{n,i}-X^{\prime}_{n,i}||_{1}$ 2000 times ($M=2000$). The initial value of $X_{0}$ satisfies equation (9) in [simconv], which is consistent with their initial values. The initial value of $X^{\prime}_{0}$ satisfies equation 7.
Figure 3 graphs the estimated upper bound on the total variation distance between the Markov chain and its corresponding stationary distribution. The differences between the simulations started off very large, but converged extremely quickly. For this reason, the y-axis is log-scaled.
Figure 3: The thick black line represents the upper bound on the Wasserstein distance on the Gibbs sampler for a variance component model (real data example 4.2). The estimate uses the mean of 2000 simulations with initial value $X_{0}$ satisfying equation (9) in [simconv] and $X_{n}=(V_{n},W_{n},\mu_{n},\vec{\theta}_{n})$. The light blue area represents the quantiles. The vertical axis is a log scale. The differences appear to be significantly right skewed.
@@
The CRN bound indicates total variation distance of less than 0.01 by iteration 554 ($||\mathcal{L}(X_{554})-\pi||_{TV}\leq 0.00529$). The DnM bound indicates total variation distance of less than 0.01 by iteration $98,750$ ($||\mathcal{L}(X_{98,750})-\pi||_{TV}\leq 0.0072$). The large disparity between the DnM and the CRN bound is consistent with the conclusions of [dmlim] that say that the convergence rate bounds based on single-step DnM tend to be overly conservative. See Figure 4 for a graphical comparison between the two bounds.
Comparison of bounds on $||\mathcal{L}(X_{n})-\pi||_{TV}$ Refer to caption
@@
Figure 4: Comparison of an upper bound on the total variation distance on the Gibbs sampler for a variance component model (real data example 4.2) using CRN simulation and a DnM bound. The CRN estimate uses the mean of 2000 simulations with initial value $X_{0}$ satisfying equation (9) in [simconv] and $X_{n}=(V_{n},W_{n},\mu_{n},\vec{\theta}_{n})$. The x-axis begins at $471$ when the CRN bound first hits below the DnM bound.
5 Gibbs sampler for a model related to James-Stein estimators
Consider a Gibbs sampler for a model related to the James-Stein estimator, which we will apply Theorem 3.3 to simulate convergence bounds on the Wasserstein distance. See [stein] for an explanation on how this model is related to James-Stein estimators. We compare our bound to the ones generated in [stein] and [sixta], which use DnM and one-shot coupling arguments, respectively, in Real Data Example 5.1.
Example 5.1 (A model related to James-Stein estimators).
$Y_{i}|\theta_{i}\sim N(\theta_{i},V)$
The Gibbs sampler for a model related to James-Stein estimators.
Algorithm 3 shows how to apply a Gibbs sampler on this example.
Initialize $(\vec{\theta}_{0},\mu_{0},A_{0})$
for $n=1,\ldots,N$ do
     For each $i\in 1,\ldots,q$, generate $\theta_{n,i}|\mu_{n-1},A_{n-1},Y\sim N_{p}(\frac{Y_{i}A_{n-1}+\mu_{n-1}V}{A_{n-1}+V},\frac{VA_{n-1}}{V+A_{n-1}})$
end for
We will bound the Wasserstein distance between a Markov chain and its corresponding stationary distribution using the Lemma 5.1.
Lemma 5.1.
Let $\{X_{n}\}_{n\geq 1}=(\vec{\theta}_{n},\mu_{n},A_{n})_{n\geq 1}$ and $\{X^{\prime}_{n}\}_{n\geq 1}=(\vec{\theta}^{\prime}_{n},\mu^{\prime}_{n},A^{\prime}_{n})_{n\geq 1}$ be two copies of the Markov chain initialized with $\vec{\theta}^{\prime}_{0}\sim N_{q}(\vec{Y},VI_{q})$, $A^{\prime}_{0}\sim IG(\alpha+(q-1)/2,\beta)$, $\mu^{\prime}_{0}\sim N(\bar{\theta}^{\prime}_{0},A^{\prime}_{0})$, and $X_{0}\perp\!\!\!\perp X^{\prime}_{0}$. Fix $p\in(1,\infty]$ and let $\pi$ be the corresponding stationary distribution. Then
$W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq\left(KE[d(X_{n},X^{\prime}_{n})^{p}]\right)^{1/p}$
Where $K=\frac{1}{\int_{B}g(\vec{\theta},\mu,A)d(\vec{\theta},\mu,A)}\frac{\Gamma(\alpha+(q-1)/2)}{\Gamma(\alpha)}\frac{1}{(2\pi\beta)^{(q-1)/2}}$ and $B\subset\mathbb{R}^{q+1}\times\mathbb{R}_{+}$ is a bounded set.
The proof is in section 8.2.
Real Data Example 5.1.
We apply our results to the batting averages ($Y$) for 18 ($q=18$) major league baseball players in 1970 [pscl]. We want to find the estimated upper bound on the Wasserstein and total variation distance for a Gibbs sampler applied to the James-Stein estimator using the baseball data. We set the priors to $\alpha=0.01,\beta=2$ and $V=Var(Y)=0.0048535$. The initial value of our Markov chains are independently drawn; $(\vec{\theta}_{0},\mu_{0},A_{0})=(100)^{q+2}$ and $(\vec{\theta}^{\prime}_{0},\mu^{\prime}_{0},A^{\prime}_{0})$ is distributed according to Lemma 5.1. Using Lemma 5.1, $K=0.0027$ and
$W_{||\cdot||_{1},1}(\mathcal{L}(\vec{\theta}_{n},\mu_{n},A_{n}),\pi)\leq 0.0027\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}||(\vec{\theta}_{n,i},\mu_{n,i},A_{n,i})-(\vec{\theta}^{\prime}_{n,i},\mu^{\prime}_{n,i},A^{\prime}_{n,i})||_{1}$
holds almost surely.
Using the CRN technique, we simulated $||(\vec{\theta}_{n,i},\mu_{n,i},A_{n,i})-(\vec{\theta}^{\prime}_{n,i},\mu^{\prime}_{n,i},A^{\prime}_{n,i})||_{1}$ one thousand times ($M=1000$) over 20 iterations ($N=20$). Figure 5 graphs the upper bound of the Wasserstein distance between $\mathcal{L}(X_{n})$ and $\pi$ on a log scale.
Upper bound on $W_{||\cdot||_{1},1}(\mathcal{L}(\theta_{n},\mu_{n},A_{n}),\pi)$ Refer to caption
Figure 5: The thick black line represents the upper bound on the Wasserstein distance for the Gibbs sampler model related to James-Stein estimators. The estimate use the mean of 1000 simulations and initial value $(\vec{\theta}_{0},\mu_{0},A_{0})=(100)^{q+2}$. The light blue area represents the simulation quantiles. The vertical axis is log scaled.
At iteration 5, the Wasserstein distance first hits below 0.01 and $W_{||\cdot||_{1},1}(\mathcal{L}(\vec{\theta}_{5},\mu_{5},A_{5}),\pi)\leq 0.00866$.
Before we bound the Markov chain and the corresponding stationary distribution in total variation we introduce Theorem 4.8 of [sixta]. Theorem 4.8 of [sixta] provides an upper bound in total variation based on the expected differences between the Markov chains through the constant $L$. This theorem will also be used to compare our bounds, which use CRN simulation, to the one-shot coupling bound.
Proposition 5.2 (Theorem 4.8 of [sixta]).
$\|\mathcal{L}(\vec{\theta}_{n},\mu_{n},A_{n})-\mathcal{L}(\vec{\theta}^{\prime}_{n},\mu^{\prime}_{n},A^{\prime}_{n})\|_{TV}$ $\leq LE[|A_{0}-A^{\prime}_{0}|]\left(\frac{1}{q}\right)^{n-1}$ (11)
where $L=\frac{(S/2)^{\frac{q-1}{2}}}{\Gamma(\frac{q-1}{2})}\left(\frac{S}{q+1}\right)^{-\frac{q-3}{2}}e^{-\frac{q+1}{2}}$ and $S=\sum_{i=1}^{q}(Y_{i}-\bar{Y})^{2}$.
Using Proposition 5.2, we get $L=0.004736$. Therefore by Lemma 4.11 of [sixta],
$\|\mathcal{L}(\vec{\theta}_{n},\mu_{n},A_{n})-\pi\|_{TV}\leq 0.000013\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}||(\vec{\theta}_{n,i},\mu_{n,i},A_{n,i})-(\vec{\theta}^{\prime}_{n,i},\mu^{\prime}_{n,i},A^{\prime}_{n,i})||_{1}$
So, by iteration 3, $\|\mathcal{L}(\vec{\theta}_{3},\mu_{3},A_{3})-\pi\|_{TV}\leq 0.0091724$.
Finally, we compare our bound to Proposition 5.2 and Theorem 4 of [stein], which use one-shot coupling and DnM arguments, respectively, to generate a theoretical bound in total variation. The methodologies are outlined in Table 2. The initial values are $(\vec{\theta}_{n},\mu_{n},A_{n})=(100)^{q+2}$.
Method Common random number Drift and minorization One shot coupling
Source Lemma 5.1 Thm 4 of [stein] Prop 5.2
Prior $A\sim IG(0.01,2)$ and $f(\mu)\propto 1$ $A\sim IG(-1,2)$ and $f(\mu)\propto 1$ $f(A)\propto 1$ and $f(\mu)\propto 1$
Bound $\|\mathcal{L}(X_{n})-\pi\|_{TV}\leq 0.000013E[|X_{n}-X^{\prime}_{n}|]$ where $X_{n},X^{\prime}_{n}$ are simulated using CRN. $\|\mathcal{L}(X_{n})-\pi\|_{TV}\leq 0.967^{n}+(1.17+\sum_{i=1}^{q}(\theta_{i}-\bar{Y})^{2})(0.935)^{n}$ $\|\mathcal{L}(X_{n})-\pi\|_{TV}\leq=0.00474E[|A_{0}-A^{\prime}_{0}|]\left(\frac{1}{18}\right)^{n-1}$
Note the minor variations in the examples, Proposition 5.2 assumes that the prior on $A$ is flat. In Theorem 4 of [stein], the prior on $A\sim IG(-1,2)$ with the intent of generating a flat distribution. In our real data example we set the prior to $A\sim IG(0.01,2)$, which could approximate a flat distribution. We will assume the minor variations are comparable.
Table 4 and figure 6 compares the bounds on total variation distance using three different methods: CRN, DnM, and one-shot coupling. The CRN and one-shot coupling bound indicates total variation distance of less than 0.01 by iteration 3. The DnM bound indicates total variation distance of less than 0.01 by iteration 249. Thus, the CRN and one-shot coupling bounds perform significantly better compared to the DnM bound.
Iteration Common Random Number Drift and Minorization One-Shot coupling
1 0.0256926 167,409.95 0.4723514
2 0.1535066 156,528.33 0.0262417
3 0.0091724 146,354.02 0.0014579
4 6.396735e-04 136,841.04 8.099304e-05
5 4.103112e-05 127,946.40 4.499613e-06
6 3.958136e-06 119,629.91 2.499785e-07
7 3.235504e-07 111,853.99 1.388769e-08
8 2.754619e-08 104,583.51 7.715386e-10
9 1.572393e-09 97,785.60 4.286326e-11
10 1.207762e-10 91,429.56 2.381292e-12
249 1.695669e-310 0.009896872 2.326398e-312
Comparison of bounds on $||\mathcal{L}(\vec{\theta}_{n},\mu_{n},A_{n})-\pi||_{TV}$ Refer to caption
The bound generated from CRN simulation and one-shot coupling are expected to be very similar since to generate the geometric convergence rate of the one-shot coupling bound (the $1/18$ in table 2) we use CRN (see [sixta, avcont, localcont]).
@@ 6 Gibbs sampler for a Bayesian linear regression model
Consider a Gibbs sampler for the Bayesian linear regression model with semi-conjugate priors (see Chapter 5 of [bayemethfin]) for which we will apply Theorem 3.3 to simulate convergence bounds on the Wasserstein distance. For this particular example, we can also provide an upper bound on the total variation distance given the Wasserstein distance using similar calculations as [sixta].
@@
Let $(\beta_{n},\sigma_{n}^{2})_{n\geq 1}$ and $(\beta^{\prime}_{n},\sigma_{n}^{{}^{\prime}2})_{n\geq 1}$ be two copies of the Markov chain with $\sigma^{{}^{\prime}2}_{0}\sim IG(\alpha^{\prime},\beta^{\prime})$ and $\beta^{\prime}_{0}\sim N_{q}(\beta_{0},\Sigma_{\beta})$ where $\alpha^{\prime}=(k+\upsilon_{0})/2$ and $\beta^{\prime}=\upsilon_{0}c_{0}^{2}/2$. Fix $p\in[1,\infty)$ and let $\pi$ be the corresponding stationary distribution. Then
$W_{d,p}(\mathcal{L}(\beta_{n},\sigma^{2}_{n}),\pi)\leq\left(\frac{(2\pi)^{q/2}\det(\Sigma_{\beta})^{1/2}}{\int_{B}g(\beta,\sigma^{2})d(\beta\times\sigma^{2})}\frac{\Gamma(\alpha^{\prime})}{\beta^{\prime\alpha^{\prime}}}E[d((\beta_{n},\sigma^{2}_{n}),(\beta_{n}^{\prime},\sigma^{2^{\prime}}_{n}))^{p}]\right)^{1/p}$
Where $B\subset\mathcal{R}^{q}\times\mathcal{R}_{+}$ is a bounded set.
@@
Convergence diagnostics for variations of the Bayesian linear regression Gibbs sampler can be found in [sixta, raj, bayesregdootika, bayesregbacklund, lineregnongauss, bayesautoreg, linregbiswas]. However, we could not find any explicit bounds on this version of the Gibbs sampler for a Bayesian linear regression (ie. the bounds were applied to more simpler examples or the analysis did not provide an explicit bound). Using example 6.1 notation, the papers ([sixta, raj, bayesregdootika]) assume that $\Sigma_{\beta}=\sigma^{2}\Sigma$ where $\Sigma$ is known. That is, the variance of $Y$ and the variance of $\beta$ are proportionally the same. Putting the results of [sixta] and [raj] together, it is shown that the geometric convergence rate that is generated when using one-shot coupling is the actual geometric convergence rate (see Theorem 3.1 of [raj] compared with Theorem 4.4 of [sixta]). The papers ([bayesregbacklund, lineregnongauss]) assume that the error of $Y|\beta,\sigma^{2}$ are non-Gaussian (to take into account outliers) and provide sufficient conditions for geometric ergodicity with respect to $k$ (dimension of $Y$) and $q$ (dimension of $\beta$). [bayesautoreg] tells us that the Bayesian regression Gibbs sampler with semi-conjugate priors is geometrically ergodic regardless of the size of $k$ compared to $q$. Finally, [linregbiswas] generates upper bounds in total variation between two copies of a linear regression Gibbs sampler model. The model in [linregbiswas] adds slightly more complexity to example 6.1 by taking $\Sigma_{\beta}$ as random. Similar to our results, the chains converge quickly.
Using Lemma 6.2 we show how an upper bound on the convergence rate in Wasserstein distance can be simulated for a real data example of the Gibbs sampler on the Bayesian regression model, Real Data Example 6.1. We further provide an estimate of the upper bound in total variation using Lemma 6.1.
@@
Suppose that we are interested in evaluating the carbohydrate consumption (Y) by age, relative weight, and protein consumption (X) for twenty male insulin-dependent diabetics. For more information on this example, see Section 6.3.1 of [glm].
We want to find the estimated upper bound on the Wasserstein and total variation distance to stationarity for a Bayesian linear regression Gibbs sampler. In this case, there are 20 observed values and 4 parameters ($k=20,q=4$). We set the priors to $\beta_{0}=\vec{0}$, $\Sigma_{\beta}=I_{4}$, $\upsilon_{0}=1$, $c_{0}^{2}=10$. The initial values of our Markov chains are $\sigma^{2}_{0}=100$, $\beta_{0}\sim N_{4}(0,I_{4})$, $\sigma^{{}^{\prime}2}_{0}\sim IG((k+\upsilon_{0})/2,\upsilon_{0}c_{0}^{2}/2)$ and $\beta^{\prime}_{0}\sim N_{4}(0,I_{4})$. We can apply Theorem 3.3 by Remark Remark, which shows that $E[||(\sigma^{2}_{n},\beta_{n})-(\sigma^{{}^{\prime}2}_{n},\beta^{\prime}_{n})||_{1}]<\infty$ for $n\in\mathbb{N}$. Thus, $K\leq 2.115$ and,
$W_{||\cdot||_{1},1}(\mathcal{L}(\sigma^{2}_{n},\beta_{n}),\pi)\leq 2.115\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}||(\sigma^{2}_{n,1},\beta_{n,i})-(\sigma^{2^{\prime}}_{n,1},\beta_{n,i})||_{1}$
@@
Using CRN simulation, we simulated $||(\sigma^{2}_{n,i},\beta_{n,i})-(\sigma^{2^{\prime}}_{n,i},\beta^{\prime}_{n,i})||_{1}$ one thousand times ($M=1000$) over 100 iterations ($N=100$). Figure 7 graphs the upper bound on the Wasserstein distance.
Upper bound on $W_{||\cdot||_{1},1}(\mathcal{L}(\sigma^{2}_{n},\beta_{n}),\pi)$ Refer to caption
Figure 7: The thick black line represents the upper bound on the Wasserstein distance on the Gibbs sampler for a Bayesian linear regression model. The estimate use the mean of 1000 simulations and initial value $\sigma^{2}_{0}=100$ and $\beta_{0}\sim N_{4}(0,I_{4})$. The light blue area represents the simulation quantiles. The vertical axis is log scaled.
At iteration 9, the Wasserstein distance first hits below 0.01 and so, $W_{||\cdot||_{1},1}(\mathcal{L}(\sigma^{2}_{9},\beta_{9}),\pi)\leq 0.0084$. In comparison the Gelman-Rubin diagnostic at iteration 9 is 1.17. For 1000 simulations, none of the simulations had autocorrelation of greater than 0.01 by the 9th iteration. This indicates that by iteration 9, the Markov chain is close to stationarity, which is consistent with our bounds.
Finally, by Lemma 6.1 the total variation distance is bounded above by $\frac{(k+\upsilon_{0})^{2}}{2\upsilon_{0}c_{0}^{2}}K=46.636$ times the expected distance between $\sigma^{2}_{n}$ and $\sigma^{2^{\prime}}_{n}$,
$||\mathcal{L}(\sigma^{2}_{n},\beta_{n})-\pi||_{TV}$ $\leq 46.636E[|\sigma^{2}_{n-1}-\sigma^{2^{\prime}}_{n-1}|]$

### v4

@@ 5 Gibbs sampler for a variance component model
We are given the yield of dyestuff from 6 batches of 5 yield measurements in grams (ie. $I=6$ and $J=5$). This data was published in Davies [13]. The data can be structured as a variance component model where $\theta_{i}$ represents the mean yield for batch $i$.
We are going to compare our Wasserstein bound to the total variation bound from [11]. [11] generates an upper bound on the total variation distance between a Markov chain and its corresponding stationary distribution using a modified version of the popular DnM condition stated in [47] (see Theorem 12 of [47] and Proposition 1 of [11]) and estimates the required DnM parameters ($\Lambda,\lambda,\epsilon$) using simulation. The hyperparameters for the DnM bound are defined in Table 7.
@@
From [11] 0.5 1 1 0 0 $10^{12}$
This paper 0.5 1 1 1 0 $10^{12}$
Table 7: Hyperparameters used in Real Data Example 5.1 and [11] (See their Section 4). The hyperparameters follow the same notation as in Equation 8.
In our comparison, we will first generate an upper bound on the Wasserstein distance between the Markov chain and the corresponding stationary distribution. Next, we will generate an upper bound on the total variation distance based on the Wasserstein distance using Proposition 24 of [40].
Proposition 24 of [40] will be used to bound the total variation distance given the Wasserstein distance. Note that in [40], the prior on $\mu\sim 1$ is flat. Since in this example $\mu\sim N(0,10^{12})$, we assume that they are the same.
Proposition 5.2 (Proposition 24 of [40]).
@@
$\mathcal{L}(X_{n})-\pi||_{TV}\leq K_{TV}W_{||\cdot||_{1},1}(\mathcal{L}(X_{n-1}),\pi)$ (11)
where $K_{TV}=(c_{1}+c_{2})J^{3/2}I$ and
@@
We bound the whole chain in Proposition 5.2 $(\vec{\theta_{n}},V_{n},W_{n},\mu_{n})$ rather than the marginal chain $(\vec{\theta_{n}},\mu_{n})$ as presented in Proposition 24 of [40] due to de-initialization properties [44]. See section 5.1 of [40] for more details.
Applying R simulation to Lemma 5.1, we calculate $K=0.00198$. The following holds,
[…]
Applying Proposition 5.2, $K_{TV}=611.4339$ and so,
[…]
The initial value of $X_{0}$ is $\vec{\theta}_{0}=$ the initial values defined in Equation 9 of [11], $V_{0}\sim IG(a1,b1)$, $W_{0}\sim IG(a2,b2)$ and $\mu_{0}=\bar{Y}$. The initial value of $X^{\prime}_{0}$ is distributed according to Equation 9.
@@
We simulated $X_{n}$ and $X^{\prime}_{n}$ 500 times ($M=500$) using CRN. Figure 5 graphs the estimated upper bound on the total variation distance between the Markov chain and its corresponding stationary distribution. Table 8 provides the total variation bound estimates and the $99\%$ confidence intervals using Lemma 3.3.
@@
The CRN bound indicates total variation distance of less than 0.01 by iteration 461 ($\mathcal{L}(X_{461})-\pi||_{TV}\leq 0.0037$). The DnM bound indicates total variation distance of less than 0.01 by iteration $98,750$ ($\mathcal{L}(X_{98,750})-\pi||_{TV}\leq 0.0072$). The two bounds are compared in Figure 5 and Table 8. The large disparity between the DnM and the CRN bound is consistent with the conclusions of [39] that say that the convergence rate bounds based on single-step DnM tend to be overly conservative. See Figure 5 for a graphical comparison between the two bounds.
Comparison of bounds on $\mathcal{L}(X_{n})-\pi||_{TV}$ Refer to caption
@@
Figure 5: Comparison of total variation distance upper bounds on the Gibbs sampler for a variance component model (Real Data Example 5.1) using CRN simulation and a DnM bound. The CRN estimate uses the mean of 500 simulations. The initial values of $X_{0}$ satisfy Equation 9 in [11].
[…]
10 3.1983e+11 0 686202.5 1210249 1.7888
50 574231.3 413.0225 919.4749 1232.997 1.7821
100 61787.44 0 301.6105 483.7103 1.7737
250 10979.74 0 127.1430 231.2048 1.7490
461 9.0992e-06 0 0.0037 0.0067 1.7149
Table 8: Total variation bound estimates and confidence intervals for the variance component model, Real Data Example 5.1. Comparison of the estimated total variation bound using CRN with the total variation bound generated from [11], which was calculated using DnM.
Finally we compare our convergence rates with informal methods in Section 8.1.3. The traceplot (Figure 9) suggests that convergence of $\mu$, $V$, and $W$ to stationarity occurs after the first few iterations (the variance of $\mu$ seems to slightly increase with $n$, however). The Gelman-Rubin statistic is evaluated at different iterations in Table 11. By iteration 250, the Gelman-Rubin statistic is close to or less than 1.05 for $V,W$, $\mu$, and $\theta_{1}$.
@@ 6 Gibbs sampler for a Bayesian linear regression model
Consider a Gibbs sampler for the Bayesian linear regression model with semi-conjugate priors (see Chapter 5 of [41]) for which we will apply Theorem 3.2 to simulate convergence bounds on the Wasserstein distance. For this particular example, we can also provide an upper bound on the total variation distance given the Wasserstein distance using similar calculations as [51].
@@
Let $(\beta_{n},\sigma_{n}^{2})_{n\geq 1}$ and $(\beta^{\prime}_{n},\sigma_{n}^{{}^{\prime}2})_{n\geq 1}$ be two copies of the Markov chain with $\sigma^{{}^{\prime}2}_{0}\sim IG(\alpha^{\prime},\beta^{\prime})$ and $\beta^{\prime}_{0}\sim N_{q}(\beta_{0},\Sigma_{\beta})$ where $\alpha^{\prime}=(k+\upsilon_{0})/2-2$ and $\beta^{\prime}=\upsilon_{0}c_{0}^{2}/2$. Fix $p\in[1,\infty)$ and let $\pi$ be the corresponding stationary distribution. Then for $r>p$
$W_{d,p}(\mathcal{L}(\beta_{n},\sigma_{n}^{2}),\pi)\leq\left(\frac{C}{L}E[d((\beta_{n},\sigma_{n}^{2}),(\beta^{\prime}_{n},\sigma_{n}^{\prime 2}))^{r}]\right)^{1/r}$
Where
  • $C=\frac{16\Gamma((k+\upsilon_{0})/2-2)(2\pi)^{q/2}\det(\Sigma_{\beta})^{1/2}}{\Gamma((k+\upsilon_{0})/2)((Y-X\hat{\beta})^{T}(Y-X\hat{\beta}))^{2}\left(\upsilon_{0}c^{2}_{0}/2\right)^{(k+\upsilon_{0})/2-2}e^{2}}$
  • $L=\int_{B}\frac{e^{-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})}}{(\upsilon_{0}c_{0}^{2}/2+(Y-X\beta)^{T}(Y-X\beta)/2)^{(k+\upsilon_{0})/2}}\mathrm{d}(\beta)$ and $B\subset\mathbb{R}^{q}\times\mathbb{R}_{+}$ is a bounded set.
@@
Convergence diagnostics for variations of the Bayesian linear regression Gibbs sampler can be found in [51, 42, 56, 4, 25, 18, 6]. However, we could not find any explicit bounds on this version of the Gibbs sampler for a Bayesian linear regression (ie. the bounds were applied to more simpler examples or the analysis did not provide an explicit bound). Using example 6.1 notation, the papers [51, 42, 56] assume that $\Sigma_{\beta}=\sigma^{2}\Sigma$ where $\Sigma$ is known. That is, the variance of $Y$ and the variance of $\beta$ are proportionally the same. Putting the results of [51] and [42] together, it is shown that the geometric convergence rate that is generated when using one-shot coupling is the actual geometric convergence rate (see Theorem 3.1 of [42] compared with Theorem 4.4 of [51]). The actual geometric convergence rate is calculated using the spectral gap in [42]. The papers ([4, 25]) assume that the error of $Y|\beta,\sigma^{2}$ are non-Gaussian (to take into account outliers) and provide sufficient conditions for geometric ergodicity with respect to $k$ (dimension of $Y$) and $q$ (dimension of $\beta$). [18] tells us that the Bayesian regression Gibbs sampler with semi-conjugate priors is geometrically ergodic regardless of the size of $k$ compared to $q$. Finally, [6] generates upper bounds in total variation between two copies of a linear regression Gibbs sampler model. The model in [6] adds slightly more complexity to example 6.1 by taking $\Sigma_{\beta}$ as random. Similar to our results, the chains converge quickly.
Using Lemma 6.2 we show how an upper bound on the convergence rate in Wasserstein distance can be simulated in Real Data Example 6.1. We further provide an estimate of the upper bound in total variation using Lemma 6.1.
@@
Suppose that we are interested in evaluating the carbohydrate consumption (Y) by age, relative weight, and protein consumption (X) for twenty male insulin-dependent diabetics. For more information on this example, see Section 6.3.1 of [16].
We want to find the estimated upper bound on the Wasserstein and total variation distance to stationarity for a Bayesian linear regression Gibbs sampler. In this case, there are 20 observed values and 4 parameters ($k=20,q=4$). We set the priors to $\beta_{0}=\vec{0}$, $\Sigma_{\beta}=I_{4}$, $\upsilon_{0}=6$, $c_{0}^{2}=140$. The initial values of our Markov chains are $\sigma^{2}_{0}=100$, $\beta_{0}\sim N_{4}(0,I_{4})$, $\sigma^{{}^{\prime}2}_{0}\sim IG(11,420)$ and $\beta^{\prime}_{0}\sim N_{4}(0,I_{4})$. We can apply Lemma 6.2 as follows, where $K=1025971$ and $r=5$.
[…]
@@
By Lemma 6.1 the total variation distance is bounded above by $\frac{(k+\upsilon_{0})^{2}}{2\upsilon_{0}c_{0}^{2}}=0.40238$ times the expected distance between $\sigma^{2}_{n}$ and $\sigma^{2}_{\infty}$ and so,
[…]
Using CRN simulation, we simulated $(\sigma^{2}_{n},\beta_{n})-(\sigma^{2^{\prime}}_{n},\beta_{n})||_{1}$ ten thousand times ($M=10000$) over 100 iterations ($N=100$). Figure 6 graphs the upper bound on the total variation distance.
Upper bound on $\mathcal{L}(\sigma^{2}_{n},\beta_{n})-\pi||_{TV}$ Refer to caption
Figure 6: Total variation distance upper bound on the Gibbs sampler for a linear regression model (Real Data Example 6.1) using CRN. The estimate uses the mean of 10000 simulations with initial values $\sigma^{2}_{0}=100$ and $\beta_{0}\sim N(0,I_{q})$. The vertical axis is log-scaled and $r=5$.
By iteration 58, the total variation distance first hits below 0.01, $\mathcal{L}(\sigma^{2}_{58},\beta_{58})-\pi||_{TV}\leq 0.0016$. In comparison, the Gelman-Rubin statistic first hits below 1.05 at iteration 49 (see Figure 10) and the traceplot suggests convergence after the first few iterations (see Figure 11).